Solve the following equations:
 3(n – 5) = 21..............

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Let us divide both the side by 3,
= (3(n - 5))/3 = 21/3
= n - 5= 7
By transposing -5 from LHS to RHS it becomes 5
= n =7 +5
= n = 12

Thanks